250 mL of argon gas is held in a flexible vessel shown above. If the pressure changes to 12.0 atm, what is the new volume of the gas at 20 °C
1
Expert's answer
2020-11-18T13:52:25-0500

pV=nRT

V=nRT/p

n (Ar) = V/22.4 = 0.25/22.4 = 0.01 mol

R = 0.082 L⋅atm K−1⋅mol−1

V (Ar)=(0.01 x 0.082 x (20+273))/12=0.022 l = 22 ml